OpenStudy (anonymous):

What is the solution for this inequality? 8x < -32 underscore <

OpenStudy (anonymous):

under score y?

OpenStudy (anonymous):

underscore < i can't do the sign

OpenStudy (anonymous):

oh write it out in english so like :eight x is less than -32 less than or equal? to some number?

OpenStudy (anonymous):

i believe you're missing a number

OpenStudy (anonymous):

or variable

OpenStudy (anonymous):

nah that's the question i'll rewerite it

OpenStudy (anonymous):

8x lessthan or equal to -32

OpenStudy (anonymous):

alright so \[8x\le-32\]

OpenStudy (anonymous):

it wants you to find x so what do you need to do get the coefficient of x to equal 1?

OpenStudy (anonymous):

yeah i think i'm not sure how to do this question though.

OpenStudy (anonymous):

\[8x \le -32 \] divide both sides by 8 = \[x \le -4\]